HUBBLE SPACE TELESCOPE DAILY REPORT #2225 PERIOD COVERED: 0000Z (UTC) 10/07/98 - 0000Z (UTC) 10/08/98 Daily Status Report as of 281/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of STIS/CCD/MA2/WF/PC-2/NIC/1/2/3 8076 (HDF-South Observations) The WF/PC-2, The Space Telescope Imaging Spectrograph (MA2 and CCD) and the Near Infrared Camera (Cameras One, Two and Three) all were used to make deep observations of the Hubble Deep Fields-South (HDFS-22333). This program is designed to provide deep observations of a high-galactic latitude field in the southern continuous viewing zone. The STIS field is centered on a bright QSO at z=2.25. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Three Sets of NIC/1/2/3 7596 (NICMOS MULTIACCUM Dark Characterization) The Near-Infrared Camera Multi-Object Spectrograph (Cameras One, Two and Three) was used to make a NICMOS MULTIACCUM dark characterization. The purpose of this proposal is to monitor the darks using three different MULTIACCUM sequences. As described in 2.1 and HSTAR 6705, the acquisition for the final iteration of this proposal defaulted to control on one FGS only. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.3 Completed STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed NIC/2 7233 (Dust Disks Around Main Sequence Stars) The Near Infrared Camera (Camera Two) was used to make observations of the dust disk around the main sequence star Alpha-Lyra (a.k.a., Vega). The observations were made with the coronagraph to minimize the effects of glare from the bright central star. A dark calibration image was also taken. As described in 2.1 and HSTAR 6705, the acquisition for this proposal defaulted to control on one FGS only. Otherwise, the observations were completed as scheduled, and no further anomalies were reported. 1.5 Completed STIS/MA1 7298 (A Snapshot Survey for UV bright Quasars at redshift below 3) The Space Telescope Imaging Spectrograph (MA1) was used to make ultraviolet snapshot observations of the bright quasar S2-1848+28 which has a redshift (Z) below 3.0. The goal is to find quasars bright enough in the ultraviolet for spectroscopic follow-up in the range OF 900-1200 Angstroms. A wave length calibration image was also taken. The observation were completed as scheduled, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 10 Successful : 10 The acquisition at 280/192733Z defaulted to fine lock on FGS-3 only when the scan step limit was exceeded on FGS-1. HSTAR 6705 was written. The proposals detailed in 1.2 and 1.4 may have been affected. Scheduled Reacquisitions : 7 Successful : 7 2.2 FHST Updates: Scheduled : 16 Successful : 16 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared three times. As described in HSTAR 6706, there was a NICMOS EDAC single bit error at 280/194156Z. Per an operations request, at 281/0330Z, the NICMOS internal log was dumped and the ground system limit was raised to 21. The HST ground system was configured to the IP network at 280/1840Z. Realtime data, at the 1Meg rate was received from 280/1947Z-1949Z and from 280/2034Z-2035Z. A TTR was written for two TDRSS events which were lost due to outage. A total of 15 minutes recoverable data was lost. Using ROP RD-7, ESTR-2 was re-conditioned at 281/0405Z and ESTR-3 was re-conditioned at 281/0807Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH